# The Tantalizing Case of the Quasar J09505128 – I. Presentation of the Data and Detailed Exploration of the Binary Supermassive Black Hole Scenario

### v3

@@ Abstract
Spectroscopic observations of the quasar J09505128 spanning 22 years reveal monotonic radial velocity variations in its broad H emission line. Moreover, the line profile becomes broader over time, necessitating careful measurements. We present robust H velocity shift measurements obtained via cross correlation, applied to both the full spectra and to isolated broad H components derived from spectral decomposition. We also examine the light curves for variability consistent with the spectroscopic trends. Using Lomb-Scargle periodogram analysis we find no significant periodic signal. We consider several interpretations for the observed changes, including a binary supermassive black hole, dust-cloud obscuration, outflows, a recoiling black hole, and a single perturbed, disk-like broad-line region. We disfavor all but the binary and perturbed broad-line region scenarios. The binary interpretation is the only one for which we can immediately compare a physical model to the available data. Thus, we incorporate radial velocity “jitter” to emulate typical quasar variability and fit the radial velocity curve with a Keplerian model to examine whether it can reproduce the observations. In this context, the available observations trace only a segment of the putative orbit. The fit yields a period of 33 years (observed frame) and an eccentricity of 0.65, with lower limits on the semi-major axis and black hole mass of pc and , respectively. Thus, J09505128 is a compelling supermassive binary candidate. The single, perturbed broad-line region interpretation remains viable but requires additional observations and modeling for further evaluation. Continued monitoring is, therefore, essential.
1 Introduction
1.1 Background and Motivation
@@
Galaxy mergers are a frequently observed phenomenon, and given that most massive galaxies have central supermassive black holes (BHs; kormendy1995inward), supermassive black hole binaries (SBHBs) are thought to be an inevitable consequence of the merger process. begelman1980massive described how a SBHB evolves in a post-galaxy merger. The evolution begins with the stellar bulges and respective BHs of the two galaxies sinking toward the center of the merger remnant. This process is initially driven by dynamical friction (chandrasekhar1943dynamical), which acts as the dominant mechanism for angular momentum loss for 109 yr for typical masses and can bring the BH pair closer. At a separation of pc, the BHs form a gravitationally bound binary (e.g., bogdanovic2022electromagnetic). Eventually, the binary “hardens” as its orbital velocity becomes comparable to the velocity dispersion of the stars. At this point, stellar scattering takes over as the dominant mechanism that further reduces the binary separation. Stellar scattering is effective so long as the number of stars with the right orbits to interact with the binary can be replenished (milosavljevic2003long; yu2005evolution). If this reservoir of stars is depleted, the binary separation may slow its evolution or stall at a separation of pc (known as the “last parsec problem”). However, several models––ranging from angular momentum loss through interactions with a circumbinary accretion disk or gaseous cloud (e.g., armitage2002accretion; escala2004role; dotti_lisa2006; cuadra2009massive) to dynamical models that consider non-spherical and anisotropic stellar distributions (e.g., merritt2004chaotic; berczik2006efficient; khan2012mergers; khan2013supermassive; holley2015galaxy)––now suggest that there are a number of sufficiently efficient mechanisms for continued orbital decay. Once the binary separation reaches 10-2 pc, angular momentum loss via gravitational waves (GWs) becomes an efficient mechanism, leading the two BHs to coalesce within a Hubble time.
@@
Unambiguous observational evidence for SBHBs has been confined to the early stages of their evolution. Galaxy mergers and widely separated (kilo-parsec and sub-kilo-parsec scale) BH pairs are observed (e.g., comerford2009inspiralling; comerford20091; liu2011dual; ellison2017discovery; hou2019active; foord2020second; goulding2019dual; shen2021dual; koss2023dual), but bound SBHBs have yet to be imaged and resolved directly. An exception is the radio-loud candidate CSO 0402+379, directly imaged through radio interferometry with a projected separation of 7 pc (rodriguez2006compact; bansal2017constraining). The recent evidence of a stochastic GW background by pulsar timing array (PTA) experiments is compatible with the low-frequency signal expected from a population of SBHBs at the last stages of their evolution (agazie2023nanograv; agazie2023binary; antoniadis2023second; reardon2023search; lee2023searching; miles2025meerkat), strongly suggesting that SBHBs must evolve to merger, emitting GWs in the process. This is promising for directly resolving nanohertz GWs (corresponding to orbital periods of tens of years) from individual SBHBs at sub-parsec separations. The upcoming Laser Interferometer Space Antenna (LISA; amaro2023astrophysics; colpi2024) is also expected to detect SBHB mergers, at higher frequencies corresponding to somewhat lower BH masses and later stages of SBHB evolution than those detected by the PTAs.
@@
SBHBs have also been invoked to explain a number of galaxy properties, including mass deficits in the cores of elliptical galaxies (e.g., milosavljevic2002cores; merritt2006cores), helical or precessing jets (e.g., roos1993precessing; romero2000preccessing), and X-shaped radio sources (e.g., merritt2002xshaped; gopalkrishna2003xshaped). Finding unambiguous SBHBs at small separations ( 0.11 pc) would greatly enhance our understanding of their evolution and their effects on their host galaxies. Thus, indirect methods have been developed to search for their observational signatures. The two most commonly used methods, which we describe further below, rely on periodic photometric variability or regular spectroscopic (radial velocity) variability of the broad emission lines. In addition to these, a few other methods have been proposed that are appropriate for SBHBs at separations of – 1 pc: peculiar profiles and relative intensities of broad emission lines (e.g., montuori11; montuori12), proper motion measurements via infrared interferometry (e.g. dexter20), periodic modulation of the polarization of scattered light (e.g. dotti22), kinematic signatures in reverberation mapping signals (e.g., wang18), or blue-red asymmetric reverberation of the broad emission lines (e.g. dotti23).
@@
The photometric method involves looking for periodic optical flux variability in the light curves of quasars (e.g., graham2015systematic; charisi2016population; liu2016systematic) under the assumption that either one or both BHs are accreting, and that the accretion rate is regulated on the binary orbital period (d2013accretion). Periodic oscillations can also occur due to relativistic Doppler boosting from an orbiting active BH (dorazio2023review). A prominent photometric SBHB candidate is OJ 287, whose long-term light curve displayed outbursts separated on a 12-year period (e.g., valtonen2008Nature; valtonen2012; komossa2023oj27_1; komossa2023oj287_2). The use of periodic photometric variations to identify binaries, especially when only a small number of apparent orbital cycles is observed, however, is subject to ambiguities (liu2018did; witt2022period; robnik2024period), since quasars exhibit stochastic variability that can appear periodic (vaughan2016false; liu2016systematic; barth_stern2018; elbadry2025arXiv).
@@
The spectroscopic method looks for the displaced peaks of broad emission lines relative to the host galaxy rest frame in quasar spectra (e.g., gaskell1983quasars; gaskell1996evidence; gaskell1996supermassive)––this is the approach used in this paper. A single-peaked broad emission line that is significantly offset is a possible signature of bulk motion of a single active BH and its surrounding gas in a binary. The possibility that the double-peaked emitters of some quasars can arise from systems with two active BHs has been tested and rejected based on the lack of radial velocity variations that are indicative of orbital motion (eracleous1997rejection; liu2016doublepeak; doan2020improved). Rather, double-peaked emission lines are best explained as originating in a disk-like broad-line region around a single supermassive black hole (see, for example, eracleous2003radioloud; eracleous2009review, for summary of arguments). While the single-peaked offset line method is viable for SBHB searches, it suffers from the ambiguity that offsets of broad emission lines can also be caused by other effects, such as outflows from the accretion disk (e.g., popovic2012super) or large, non-axisymmetric perturbations of the broad-line region (BLR; e.g., gezari2007long; lewis2010long; storchi2017double; schimoia2017evolution). It is therefore necessary to have multiple observations over extended periods of time to evaluate SBHB candidates identified by this method.
@@
In order to select SBHB candidates based on their displaced broad emission lines, one approach is to search for 1000 km s-1 velocity offsets of the broad H emission lines relative to the narrow emission lines in the spectra of quasars (gaskell1983quasars; tsalmantza2011systematic; eracleous2012large; decarli2013nature; liu2014constraining). Using this method, eracleous2012large selected 88 SBHB candidates which have since undergone extensive studies to characterize their nature such as flux variability, radial velocity changes, and implied population properties (runnoe2015large; runnoe2017large; pflueger2018likelihood; nguyen2020pulsar). runnoe2017large used observations through 2011 and found 29 candidates that displayed steady, and in some cases, statistically significant radial velocity variations. Among these, three objects displayed monotonic changes consistent with the expected behavior of SBHBs. These objects became prime targets for further follow-up observations. After obtaining new observations over several years using ground-based telescopes, we identify J095036.75512838.1 (hereafter J0950) as a promising SBHB candidate.
@@
1.2 Possible Interpretations of a Shifted Broad Emission Line and Scope of this Paper
@@
Here we explore several interpretations for the apparent broad emission line shift of J0950. These include: (a) a binary supermassive black hole interpretation, which is the hypothesis originally considered in previous works on this object (Section 1.1), (b) a dust cloud outside the BLR obscuring parts of it as it moves across the line of sight, thereby producing asymmetric broad emission lines that mimic radial velocity shifts; (c) outflows, where the delays in the arrival times of ejecta moving towards and away from the observer result in apparent blueshifts and redshifts of the broad emission lines across different epochs; (d) a recoiling BH that has been kicked out of, and subsequently falls back into, the gravitational potential well of its host galaxy; and (e) a perturbation in a disk-like BLR that can produce a broad emission line peak that shifts in velocity as the perturbation precesses. We revisit these interpretations for further consideration in the Section 5.2.
@@
The goal of this paper is to present the available data and a thorough study of the variability of the broad H line, and to explore the SBHB interpretation in detail. It is the first in a series of papers that will investigate possible scenarios for the behavior of this object through observational tests and/or modeling. In Section 2, we present the properties of the target and available spectroscopic and photometric data. Section 3 describes the methods for the velocity measurements, simulations of the observed profile change of the broad H emission line, and the relative velocities resulting from the H shift measurements. In Section 4, we present the analysis of the light curves. We assess the different interpretations of the observed behavior in J0950’s spectra and provide a detailed discussion of the SBHB interpretation in Section 5. Section 6 summarizes our results and notes future work. We assume a standard cosmology with , , and (bennett_cosmology2014) throughout this work.111We calculated luminosity distance using Ned Wright’s cosmology calculator (wright2006cosmology), available at https://www.astro.ucla.edu/~wright/CosmoCalc.html
@@
2 Properties of the Target and Available Data
2.1 Properties of the Target
J0950 is one of 88 SBHB candidates selected by eracleous2012large from a sample of 15,900 Sloan Digital Sky Survey (SDSS) quasars. The candidates were identified via velocity offsets of their broad H emission lines from the redshift defined by their [O iii] 4959,5007 narrow lines. Its redshift, measured from the rest wavelength of [O iii] 5007, is , which implies a luminosity distance of 1067.9 Mpc. Its absolute and apparent V-band magnitudes are 21.75 and 18.35 respectively. The short-term flux variability properties of the continuum and broad H emission of J0950 are much like those observed in normal quasars in the same redshift and luminosity range (runnoe2015large), making it appear to be a normal quasar apart from its offset broad-line.
2.2 Available Spectra
The current spectroscopic dataset for J0950 consists of spectra taken with the SDSS 2.5m telescope, the Palomar Hale 5m Telescope (Pal), the Keck 10m telescope, and the Hobby-Eberly 11m Telescope (HET) on the dates shown in Table 1 and Figure 1. The calibration process for the early spectra is described in Section 5.1 of eracleous2012large and summarized in Section 2.1 of runnoe2015large. The recent spectra from the HET were taken with the second-generation Low-Resolution Spectrograph (LRS2; lrs2_2016SPIE.9908E..4CC), which is fed by a fiber bundle. The LRS2 spectra were reduced and calibrated with the Panacea222https://github.com/grzeimann/Panacea package, written by Greg Zeimann. The reduction consists of bias and dark subtraction, fiber tracing and extraction as well as wavelength calibration, fiber-to-fiber normalization, source detection and extraction, and flux calibration. Starting with those spectra, we carried out the following corrections:
@@ 2.2 Available Spectra
The flux density was corrected for Galactic interstellar extinction using the schlegel1998maps dust maps333The dust map was accessed using sfdmap (codebase: https://github.com/kbarbary/sfdmap) and extinction (barbary_k_2016) python packages. recalibrated by schlafly2011measuring, and the extinction law of fitzpatrick1999correcting.
The wavelength scale of each spectrum was converted from air to vacuum using the formula provided by key, which is an inverse transformation of the vacuum-to-air conversion given by morton2000atomic.
To verify the absolute wavelength calibration (hence the velocity stability) of the spectra, we used the [O iii] 4959,5007 doublet, which originates on large scales and serves as an internal wavelength standard. We found that, relative to the 2002 SDSS spectrum the subsequent spectra had shifts between and km s-1, as determined using the cross-correlation method described in Section 3. We corrected these shifts to adjust the spectral alignment. The residual uncertainty after the alignment process corresponds to km s-1, which is set by how well we can measure the shift in the [O iii] doublet via cross-correlation.
2.3 Available Light Curves

### v4

@@ Abstract
Spectroscopic observations of the quasar J09505128 spanning 22 years reveal monotonic radial velocity variations in its broad H emission line. Moreover, the line profile becomes broader over time, necessitating careful measurements. We present robust H velocity shift measurements obtained via cross correlation, applied to both the full spectra and to isolated broad H components derived from spectral decomposition. We also examine the light curves for variability consistent with the spectroscopic trends. Using Lomb-Scargle periodogram analysis we find no significant periodic signal. We consider several interpretations for the observed changes, including a binary supermassive black hole, dust-cloud obscuration, outflows, a recoiling black hole, and a single perturbed, disk-like broad-line region. We deem the binary and perturbed broad-line region scenarios to be physically plausible. The binary interpretation is the only one for which we can immediately compare a physical model to the available data. Thus, we incorporate radial velocity “jitter” to emulate typical quasar variability and fit the radial velocity curve with a Keplerian model to examine whether it can reproduce the observations. In this context, the available observations trace only a segment of the putative orbit. The fit yields a period of 33 years (observed frame) and an eccentricity of 0.65, with lower limits on the semi-major axis and black hole mass of pc and , respectively. Thus, J09505128 is a binary candidate deserving further study. The single, perturbed broad-line region interpretation remains viable but requires additional observations and modeling for further evaluation. Continued monitoring is, therefore, essential.
I Introduction
I.1 Background and Motivation
@@
Galaxy mergers are a frequently observed phenomenon, and given that most massive galaxies have central supermassive black holes (BHs; Kormendy and Richstone 1995), supermassive black hole binaries (SBHBs) are thought to be an inevitable consequence of the merger process. Begelman et al. (1980) described how a SBHB evolves in a post-galaxy merger. The evolution begins with the stellar bulges and respective BHs of the two galaxies sinking toward the center of the merger remnant. This process is initially driven by dynamical friction (Chandrasekhar 1943), which acts as the dominant mechanism for angular momentum loss for 109 yr for typical masses and can bring the BH pair closer. At a separation of pc, the BHs form a gravitationally bound binary (e.g., Bogdanović et al. 2022). Eventually, the binary “hardens” as its orbital velocity becomes comparable to the velocity dispersion of the stars. At this point, stellar scattering takes over as the dominant mechanism that further reduces the binary separation. Stellar scattering is effective so long as the number of stars with the right orbits to interact with the binary can be replenished (Milosavljević and Merritt 2003; Yu et al. 2005). If this reservoir of stars is depleted, the binary separation may slow its evolution or stall at a separation of pc (known as the “last parsec problem”). However, several models––ranging from angular momentum loss through interactions with a circumbinary accretion disk or gaseous cloud (e.g., Armitage and Natarajan 2002; Escala et al. 2004; Dotti et al. 2006; Cuadra et al. 2009) to dynamical models that consider non-spherical and anisotropic stellar distributions (e.g., Merritt and Poon 2004; Berczik et al. 2006; Khan et al. 2012, 2013; Holley-Bockelmann and Khan 2015)––now suggest that there are a number of sufficiently efficient mechanisms for continued orbital decay. Once the binary separation reaches 10-2 pc, angular momentum loss via gravitational waves (GWs) becomes an efficient mechanism, leading the two BHs to coalesce within a Hubble time.
@@
Unambiguous observational evidence for SBHBs has been confined to the early stages of their evolution. Galaxy mergers and widely separated (kilo-parsec and sub-kilo-parsec scale) BH pairs are observed (e.g., Comerford et al. 2009a, b; Liu et al. 2011; Ellison et al. 2017; Hou et al. 2019; Foord et al. 2020; Goulding et al. 2019; Shen et al. 2021; Koss et al. 2023), but bound SBHBs have yet to be imaged and resolved directly. An exception is the radio-loud candidate CSO 0402+379, directly imaged through radio interferometry with a projected separation of 7 pc (Rodriguez et al. 2006; Bansal et al. 2017). The recent evidence of a stochastic GW background by pulsar timing array (PTA) experiments is compatible with the low-frequency signal expected from a population of SBHBs at the last stages of their evolution (Agazie et al. 2023a, b; EPTA Collaboration et al. 2023; Reardon et al. 2023; Xu et al. 2023; Miles et al. 2025), strongly suggesting that SBHBs must evolve to merger, emitting GWs in the process. This is promising for directly resolving nanohertz GWs (corresponding to orbital periods of tens of years) from individual SBHBs at sub-parsec separations. The upcoming Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al. 2023; Colpi et al. 2024) is also expected to detect SBHB mergers, at higher frequencies corresponding to somewhat lower BH masses and later stages of SBHB evolution than those detected by the PTAs.
@@
SBHBs have also been invoked to explain a number of galaxy properties, including mass deficits in the cores of elliptical galaxies (e.g., Milosavljević et al. 2002; Merritt 2006), helical or precessing jets (e.g., Roos et al. 1993; Romero et al. 2000), and X-shaped radio sources (e.g., Merritt and Ekers 2002; Gopal-Krishna et al. 2003). Finding unambiguous SBHBs at small separations ( 0.11 pc) would greatly enhance our understanding of their evolution and their effects on their host galaxies. Thus, indirect methods have been developed to search for their observational signatures. The two most commonly used methods, which we describe further below, rely on periodic photometric variability or regular spectroscopic (radial velocity) variability of the broad emission lines. In addition to these, a few other methods have been proposed that are appropriate for SBHBs at separations of – 1 pc: peculiar profiles and relative intensities of broad emission lines (e.g., Montuori et al., 2011, 2012), proper motion measurements via infrared interferometry (e.g. Dexter et al., 2020), periodic modulation of the polarization of scattered light (e.g. Dotti et al., 2022), kinematic signatures in reverberation mapping signals (e.g., Wang et al., 2018), or blue-red asymmetric reverberation of the broad emission lines (e.g. Dotti et al., 2023).
@@
The photometric method involves looking for periodic optical flux variability in the light curves of quasars (e.g., Graham et al. 2015; Charisi et al. 2016; Liu et al. 2016b) under the assumption that either one or both BHs are accreting, and that the accretion rate is regulated on the binary orbital period (D’Orazio et al. 2013). Periodic oscillations can also occur due to relativistic Doppler boosting from an orbiting active BH (D’Orazio and Charisi 2023). A prominent photometric SBHB candidate is OJ 287, whose long-term light curve displayed outbursts separated on a 12-year period (e.g., Valtonen et al. 2008, 2012; Komossa et al. 2023b, a). The use of periodic photometric variations to identify binaries, especially when only a small number of apparent orbital cycles is observed, however, is subject to ambiguities (Liu et al. 2018; Witt et al. 2022; Robnik et al. 2024), since quasars exhibit stochastic variability that can appear periodic (Vaughan et al. 2016; Liu et al. 2016b; Barth and Stern 2018; El-Badry et al. 2025).
@@
The spectroscopic method looks for the displaced peaks of broad emission lines relative to the host galaxy rest frame in quasar spectra (e.g., Gaskell 1983, 1996b, 1996a)––this is the approach used in this paper. A single-peaked broad emission line that is significantly offset is a possible signature of bulk motion of a single active BH and its surrounding gas in a binary. The possibility that the double-peaked emitters of some quasars can arise from systems with two active BHs has been tested and rejected based on the lack of radial velocity variations that are indicative of orbital motion (Eracleous et al. 1997; Liu et al. 2016a; Doan et al. 2020). Rather, double-peaked emission lines are best explained as originating in a disk-like broad-line region around a single supermassive black hole (see, for example, Eracleous and Halpern, 2003; Eracleous et al., 2009, for summary of arguments). While the single-peaked offset line method is viable for SBHB searches, it suffers from the ambiguity that offsets of broad emission lines can also be caused by other effects, such as outflows from the accretion disk (e.g., Popović 2012) or large, non-axisymmetric perturbations of the broad-line region (BLR; e.g., Gezari et al. 2007; Lewis et al. 2010; Storchi-Bergmann et al. 2017; Schimoia et al. 2017). It is therefore necessary to have multiple observations over extended periods of time to evaluate SBHB candidates identified by this method.
@@
In order to select SBHB candidates based on their displaced broad emission lines, one approach is to search for 1000 km s-1 velocity offsets of the broad H emission lines relative to the narrow emission lines in the spectra of quasars (Gaskell 1983; Tsalmantza et al. 2011; Eracleous et al. 2012; Decarli et al. 2013; Liu et al. 2014). Using this method, Eracleous et al. (2012) selected 88 SBHB candidates which have since undergone extensive studies to characterize their nature such as flux variability, radial velocity changes, and implied population properties (Runnoe et al. 2015, 2017; Pflueger et al. 2018; Nguyen et al. 2020). Runnoe et al. (2017) used observations through 2011 and found 29 candidates that displayed steady, and in some cases, statistically significant radial velocity variations. Among these, three objects displayed monotonic changes consistent with the expected behavior of SBHBs. These objects became prime targets for further follow-up observations. After obtaining new observations over several years using ground-based telescopes, we identify J095036.75512838.1 (hereafter J0950) as a promising SBHB candidate.
@@
I.2 Possible Interpretations of a Shifted Broad Emission Line and Scope of this Paper
@@
Here we explore several interpretations for the apparent broad emission line shift of J0950. These include: (a) a binary supermassive black hole interpretation, which is the hypothesis originally considered in previous works on this object (Section I.1), (b) a dust cloud outside the BLR obscuring parts of it as it moves across the line of sight, thereby producing asymmetric broad emission lines that mimic radial velocity shifts; (c) outflows, where the delays in the arrival times of ejecta moving towards and away from the observer result in apparent blueshifts and redshifts of the broad emission lines across different epochs; (d) a recoiling BH that has been kicked out of, and subsequently falls back into, the gravitational potential well of its host galaxy; and (e) a perturbation in a disk-like BLR that can produce a broad emission line peak that shifts in velocity as the perturbation precesses. We revisit these interpretations for further consideration in the Section V.2.
@@
The goal of this paper is to present the available data and a thorough study of the variability of the broad H line, and to explore the SBHB interpretation in detail. It is the first in a series of papers that will investigate possible scenarios for the behavior of this object through observational tests and/or modeling. In Section II, we present the properties of the target and available spectroscopic and photometric data. Section III describes the methods for the velocity measurements, simulations of the observed profile change of the broad H emission line, and the relative velocities resulting from the H shift measurements. In Section IV, we present the analysis of the light curves. We assess the different interpretations of the observed behavior in J0950’s spectra and provide a detailed discussion of the SBHB interpretation in Section V. Section VI summarizes our results and notes future work. We assume a standard cosmology with , , and (Bennett et al., 2014) throughout this work.111We calculated luminosity distance using Ned Wright’s cosmology calculator (Wright, 2006), available at https://www.astro.ucla.edu/~wright/CosmoCalc.html
@@
II Properties of the Target and Available Data
II.1 Properties of the Target
J0950 is one of 88 SBHB candidates selected by Eracleous et al. (2012) from a sample of 15,900 Sloan Digital Sky Survey (SDSS) quasars. The candidates were identified via velocity offsets of their broad H emission lines from the redshift defined by their [O iii] 4959,5007 narrow lines. Its redshift, measured from the rest wavelength of [O iii] 5007, is , which implies a luminosity distance of 1067.9 Mpc. Its absolute and apparent V-band magnitudes are 21.75 and 18.35 respectively. The short-term flux variability properties of the continuum and broad H emission of J0950 are much like those observed in normal quasars in the same redshift and luminosity range (Runnoe et al. 2015), making it appear to be a normal quasar apart from its offset broad-line.
II.2 Available Spectra
The current spectroscopic dataset for J0950 consists of spectra taken with the SDSS 2.5m telescope, the Palomar Hale 5m Telescope (Pal), the Keck 10m telescope, and the Hobby-Eberly 11m Telescope (HET) on the dates shown in Table 1 and Figure 1. The calibration process for the early spectra is described in Section 5.1 of Eracleous et al. (2012) and summarized in Section 2.1 of Runnoe et al. (2015). The recent spectra from the HET were taken with the second-generation Low-Resolution Spectrograph (LRS2; Chonis et al. 2016), which is fed by a fiber bundle. The LRS2 spectra were reduced and calibrated with the Panacea222https://github.com/grzeimann/Panacea package, written by Greg Zeimann. The reduction consists of bias and dark subtraction, fiber tracing and extraction as well as wavelength calibration, fiber-to-fiber normalization, source detection and extraction, and flux calibration. Starting with those spectra, we carried out the following corrections:
@@ II.2 Available Spectra
The flux density was corrected for Galactic interstellar extinction using the Schlegel et al. (1998) dust maps333The dust map was accessed using sfdmap (codebase: https://github.com/kbarbary/sfdmap) and extinction (Barbary 2016) python packages. recalibrated by Schlafly and Finkbeiner (2011), and the extinction law of Fitzpatrick (1999).
To verify the absolute wavelength calibration (hence the velocity stability) of the spectra, we used the [O iii] 4959,5007 doublet, which originates on large scales and serves as an internal wavelength standard. We found that, relative to the 2002 SDSS spectrum the subsequent spectra had shifts between and km s-1, as determined using the cross-correlation method described in Section III. We corrected these shifts to adjust the spectral alignment. The residual uncertainty after the alignment process corresponds to km s-1, which is set by how well we can measure the shift in the [O iii] doublet via cross-correlation.
II.3 Available Light Curves